OpenStudy (otherworldly):

Which statement is true about the difference 2 square root of 7 − square root of 28? It is rational and equal to −2. It is rational and equal to 0. It is irrational and equal to −2square root of 7. It is irrational and equal to square root of 7. @DanJS

OpenStudy (danjs):

this? \[\large 2\sqrt{7}-\sqrt{28}\]

OpenStudy (otherworldly):

yes

OpenStudy (danjs):

you cant simplify the first 2root7, the root28 you can though...

OpenStudy (otherworldly):

yes is it square root of 7 x square root of 2 and square root of 2 then its 4square root of 7 ???

OpenStudy (danjs):

28=4*7 root 4 is 2, so the root28 goes to 2root7 \[\sqrt{28}=\sqrt{4*7}=\sqrt{4}\sqrt{7}=2\sqrt{7}\] \[2\sqrt{7}-\sqrt{28}=2\sqrt{7}-2\sqrt{7}=0\]

OpenStudy (otherworldly):

oh ok

OpenStudy (otherworldly):

i must have made a mistake somewhere ....thank you :)

OpenStudy (danjs):

welcome